OpenStudy (anonymous):

x^3-4x+1=0 Solve Using Rational Root Theorem

OpenStudy (anonymous):

\[x^3-4x+1=0\]

OpenStudy (p0sitr0n):

you need to find integers such that a|coefficient of leading x and b|constant term, then a/b will give possible rationnal roots, but you still have to check them. Keep in mind that if m|n,n|m then m=+-n so a=+-1 b=+-1 thus possible roots are +-1.Lets check -1+4+1=4 1-4+1=-2 We see that neither satisfies the polynomial thus it does not have rationnal roots NOTE: it does not mean the polynomial does not have roots at all, just no rationnal ones